i. What is money?

ii.   In commercial bank money creation process, what is cash reserve? 

If GHC10,000 is deposited in a commercial bank and the cash reserve ratio is 10% of deposits, 

How much money can the bank create?      

What is the difference in the amount of money created if the cash reserve ratio changes from 10% to 5%?  

What is the level of the money multiplier?      



1
Expert's answer
2021-08-26T11:43:23-0400

i.

Money entails any asset that can be utilized to buy products.

ii.

Cash reserve refers to the amount of money the central bank requires commercial banks to keep aside and not loan it to clients.

If the cash reserve ratio is 10%

  • Money multiplier = 1 / cash reserve ratio
  • Cash reserve ratio = 10% (0.1)

Therefore;

  • Money multiplier = 1 / 0.1
  • Money multiplier = 10
  • Money the bank create = money multiplier * money deposited
  • Money the bank create = 10 * 10,000 = 100,000

If the cash reserve ratio changes from 10% to 5%

  • Money multiplier = 1 / 0.05
  • Money multiplier = 20
  • Money the bank create = money multiplier * money deposited
  • Money the bank create = 20 * 10,000 = 200,000